# Token Perturbation Guidance for Diffusion Models

### v1

@@ Abstract
Classifier-free guidance (CFG) has become an essential component of modern diffusion models to enhance both generation quality and alignment with input conditions. However, CFG requires specific training procedures and is limited to conditional generation. To address these limitations, we propose Token Perturbation Guidance (TPG), a novel method that applies perturbation matrices directly to intermediate token representations within the diffusion network. TPG employs a norm-preserving shuffling operation to provide effective and stable guidance signals that improve generation quality without architectural changes. As a result, TPG is training-free and agnostic to input conditions, making it readily applicable to both conditional and unconditional generation. We further analyze the guidance term provided by TPG and show that its effect on sampling more closely resembles CFG compared to existing training-free guidance techniques. Extensive experiments on SDXL and Stable Diffusion 2.1 show that TPG achieves nearly a 2$\times$ improvement in FID for unconditional generation over the SDXL baseline, while closely matching CFG in prompt alignment. These results establish TPG as a general, condition-agnostic guidance method that brings CFG-like benefits to a broader class of diffusion models. The code is available at https://github.com/TaatiTeam/Token-Perturbation-Guidance
@@ 1 Introduction
Diffusion models ho2020denoising ; song2020denoising ; song2020score have emerged as the main methodology behind many successful generative models for images esser2024scaling ; podell2023sdxl , videos blattmann2023stable ; bar2024lumiere ; yang2024cogvideox ; yu2024efficient ; tian2024videotetris , audio tian2025audiox ; xue2024auffusion , and 3D objects cui2024neusdfusion ; liu2024pi3d ; yan2024dreamdissector . Despite their theoretical capacity to produce high-fidelity data, unguided diffusion models often suffer from poor sample quality, manifesting as visual artifacts, lack of semantic consistency, and insufficient sharp details (karras2024guiding, ). To mitigate these issues, classifier-free guidance (CFG) ho2022classifier has become the de facto approach to steer the generation process toward higher quality and more semantically aligned outputs. However, CFG is inherently limited to conditional generation and requires a specific training strategy that randomly replaces the input condition with a null condition.
@@
In this paper, we revisit existing attention-based guidance techniques and aim to bridge the gap between the effectiveness of CFG and that of training-free, condition-agnostic methods. We observe that CFG effectively recovers global structure and coarse details during the early denoising steps, whereas existing training-free methods tend to produce over-smoothed results at the same stage (see highlighted regions in Figure 1). This is problematic because early denoising steps are critical for both image quality and prompt alignment, as they establish global structure, major shapes, and coarse semantics before the network begins refining fine details wang2023diffusion ; dhariwal2021diffusion . If the model fails to capture correct semantics, object placement, or overall composition at this stage, it may never fully recover from that high-level mismatch in later refinements. This lack of sufficient early-step guidance likely explains why existing methods often yield only marginal improvements in prompt alignment and generation quality compared to CFG.
@@
Motivated by these insights, we introduce Token Perturbation Guidance (TPG), a novel method that directly perturbs intermediate token representations within the diffusion network, without requiring additional training or architectural changes. TPG employs token shuffling as the core operation to provide effective guidance signals. Specifically, token shuffling is (i) linear, (ii) preserves token norms, and (iii) disrupts local structure while maintaining global statistics. As shown in Figure 1, TPG exhibits behavior similar to CFG and, compared to other training-free methods, more faithfully recovers both global structure and fine details at early denoising stages.
We evaluate the effectiveness of TPG on both conditional and unconditional image generation using SDXL podell2023sdxl and Stable Diffusion 2.1 rombach2022high . Our results show that TPG achieves nearly a 2$\times$ improvement in FID for unconditional generation compared to the SDXL baseline and also significantly outperforms existing perturbation-based guidance techniques. Furthermore, we observe that TPG closely mirrors CFG in terms of alignment and its effect on the sampling process, i.e., positively aligning with ground-truth noise in low-frequency bands, remaining largely orthogonal at other frequencies, and following a similar norm profile throughout denoising. These findings establish TPG as a general, condition-agnostic guidance method that extends CFG-like benefits to a broader class of diffusion models, including those for unconditional generation.
@@ 2 Related work
Score-based diffusion models sohl2015deep ; ho2020denoising ; song2019generative ; song2020score reverse a forward noising process by learning the score function, the gradient of the log data density, at multiple noise levels to progressively transform pure Gaussian noise into realistic samples (song2019generative, ). This principled estimator of the data distribution has outperformed previous generative modeling methods in both fidelity and mode coverage dhariwal2021diffusion . Since the score represents an explicit gradient field, the sampling trajectory can be guided by incorporating auxiliary gradients, leading to powerful guidance techniques such as classifier guidance (CG) dhariwal2021diffusion and classifier-free guidance (CFG) ho2022classifier . These guidance methods significantly enhance image quality and prompt alignment, albeit at the cost of oversaturation (sadat2025eliminating, ) and reduced diversity (ho2022classifier, ; sadat2024cads, ).
@@ 3 Background
where $\nabla_{\mathbf{x}}\log p_{t}({{\mathbf{x}}})$ is the score function, representing the gradient of the log-density of the noisy data at time $t$. This gradient indicates the direction in which a sample should be updated to increase its likelihood w.r.t. the noisy data distribution. The unkown score function is typically approximated by a neural network $\mathbf{s}_{\theta}(\mathbf{x},t)$, trained using denoising score matching vincent2011connection . Given a noisy sample ${\mathbf{x}}_{t}$ at time $t$, the network predicts the score of the marginal distribution $p_{t}$. Training is performed by minimizing a weighted denoising score matching loss:
@@ 4 Token perturbation guidance
  • Norm preservation: The shuffling matrix is orthonormal, meaning it satisfies ${\boldsymbol{S}}^{\top}{\boldsymbol{S}}={\boldsymbol{I}}$ and acts as a rigid rotation or permutation in the embedding space. This property guarantees that the preturbation preserves token norms, i.e., we have $||{\boldsymbol{S}}{\boldsymbol{H}}||_{2}=||{\boldsymbol{H}}||_{2}$. As a result, the magnitude of the feature representations remains unchanged, which helps maintain their statistical properties and prevents internal covariate shift ioffe2015batch .
@@ 5 Comparing the behavior of TPG with other guidance methods
As shown in Figure 2, TPG and CFG produce guidance vectors that are nearly orthogonal to the ground-truth noise throughout the trajectory, as indicated by cosine values close to zero. In contrast, PAG and SEG exhibit strong negative alignment during the middle steps, suggesting that they temporarily oppose the denoising direction. Figure 2 (b) further shows that, compared to SEG and PAG, TPG more closely mirrors the behavior of CFG in terms of alignment between the predicted and ground-truth noise. Additionally, Figure 2 (c) shows that TPG and CFG exhibit similar guidance magnitudes across all denoising steps, while SEG and PAG show substantially lower norms. This indicates that the update terms in CFG and TPG are more influential throughout the sampling process compared to that in PAG and SEG.
Figure 3 complements our step-wise analysis by illustrating how each guidance method behaves in the frequency domain. TPG and CFG remain almost perfectly orthogonal to the ground-truth noise across all frequencies and time steps, except for a slight positive tilt in the lowest frequency bands. SEG, however, exhibits a clear negative stripe at medium frequencies during intermediate steps, confirming that its correction momentarily opposes the desired direction. The norm heatmaps further reveal that CFG and TPG inject a strong low-frequency signal in the early steps, while high-frequency modifications primarily occur in the later denoising steps. In contrast, SEG operates with less energy and displays a markedly different norm pattern when modifying high-frequency content, indicating a relatively weaker approach to detail refinement. This behavior aligns with Figure 1, where the reduced energy across frequency bands leads to overly smooth generations in the initial steps.
@@ 6 Experiments
All experiments are conducted on RTX6000 GPUs via official implementations and pre-trained checkpoints provided by publicly available repositories. We build upon the current open-source state-of-the-art Stable Diffusion XL (SDXL) podell2023sdxl as our primary baseline, and include the Stable Diffusion 2.1 (SD 2.1) CompVis2022StableDiffusionv12 to show the generality of our method and analysis. Moreover, TPG is easy to implement, and it can be added into existing diffusion models as a plug-and-play module with just a few lines of additional code. In all experiments, the compared methods are evaluated using their original configurations and default guidance scales. The guidance scale for TPG is fixed at 3.0.
@@ 6.1 Quantitative results
Table 1 presents a quantitative comparison of TPG against vanilla SDXL, PAG, SEG, and CFG for both unconditional and conditional image generation tasks. In unconditional generation, TPG outperforms all compared methods by achieving notably lower FID and sFID scores (69.31 and 44.18, respectively), indicating better overall image quality and diversity. Additionally, TPG achieves the highest Inception Score (17.99) while maintaining a competitive Aesthetic Score, further reflecting the perceptual quality of TPG outputs. In conditional generation, CFG achieves the best results overall, with TPG closely following and consistently outperforming vanilla SDXL, PAG, and SEG in FID, sFID, and CLIP Score metrics.
Table 2 further compares TPG against vanilla Stable Diffusion 2.1, PAG, and SEG in conditional image generation. TPG consistently achieves the best performance, significantly outperforming the other methods with the lowest FID score (16.69) and highest Inception Score (36.28). It also maintains competitive results in terms of Aesthetic Score (5.97) and CLIP Score (29.30), highlighting TPG’s strength in producing high-quality and semantically aligned images.
@@ 6.2 Qualitative results
Figure 4 highlights the differences in unconditional generations produced by various guidance methods. Vanilla SDXL, PAG, and SEG often generate abstract or repetitive textures lacking clear semantic structure, while TPG consistently produces well-structured and realistic scenes. Across various initial seeds, TPG is less prone to generating abstract patterns and more likely to form coherent spatial layouts with identifiable objects. Moreover, visual artifacts such as unnatural textures or distortions are more noticeable in outputs from baseline methods, whereas TPG reduces such artifacts, resulting in cleaner and more realistic generations.
We also show a qualitative comparison of conditional image generation across different guidance methods in Figure 5. As can be seen, CFG and TPG consistently produce the highest-quality images, with sharp details and strong alignment to the text prompts. In contrast, other baselines such as SEG and PAG often generate outputs that deviate from the given condition, leading to less faithful semantic content. Furthermore, visual artifacts—such as distorted shapes or inconsistent textures—are more frequently observed in SEG and PAG outputs, while TPG exhibits improved robustness, generating cleaner and more semantically accurate images with fewer artifacts.
@@
Table 2: Quantitative comparison of TPG with vanilla Stable Diffusion 2.1 rombach2022high , PAG ahn2024self , and SEG hong2024smoothed for conditional generation. TPG outperforms other baselines in all evaluated metrics.
@@
[…]
6.3 Ablating other norm-preserving perturbation methods
To assess the importance of shuffling during guidance, Table 3 reports the results of various norm-preserving perturbation methods compared to the vanilla baseline. In addition to the shuffling strategy used in TPG, we evaluated Sign Flip, Hadamard, and Haar transforms, all described in Appendix A. While these three alternatives provide slight improvements in generation quality over the baseline, their gains are modest. In contrast, shuffling yields a substantially larger improvement in both FID and Inception Score, indicating significantly better image quality and diversity. Although all perturbations are orthogonal and preserve token norms, they influence the features differently. Shuffling randomly reorders tokens, disrupting local patterns while preserving recoverable global structure, which facilitates stronger guidance during inference. By comparison, Hadamard and Haar transformations mix all tokens together, potentially distorting useful information and weakening the guidance signal. Sign Flip merely alters the sign of each token, which may not offer a strong enough signal to steer samples effectively toward desirable regions of the data distribution.
@@ 7 Conclusion and discussion
Despite these strengths, as with CFG itself, TPG still requires two forward passes through the diffusion network, leading to increased sampling time compared to the unguided case. Additionally, although TPG significantly improves quality in most situations, the guidance term may remain limited in extreme out-of-distribution scenarios that are not captured by the learned distribution of the base model. We consider addressing these limitations an interesting direction for future research.
@@ Appendix A Orthogonal token perturbation matrix designs
TPG directly applies structured perturbations to the intermediate token embeddings within the denoiser during inference. Specifically, consider the intermediate hidden-state activations of the denoiser at a given layer, represented as a tensor ${\boldsymbol{H}}\in\mathbb{R}^{B\times N\times C}$, where $B$ denotes the batch size, $N$ the number of tokens, and $C$ the dimension of each token’s feature vector. At each denoiser layer $k$ and diffusion time step $i$, we apply the perturbation by multiplying the token-embedding with an orthogonal (or approximately orthogonal) matrix ${\boldsymbol{P}}_{k,i}\;\in\;\mathbb{R}^{N\times N}$ along the token dimension:
@@
  • Walsh–Hadamard Transform (WHT). The WHT uses a normalized Hadamard matrix ${\boldsymbol{W}}\in\mathbb{R}^{N\times N}$, whose entries are ±$\pm 1/\sqrt{N}$ and which satisfies ${\boldsymbol{W}}^{\top}{\boldsymbol{W}}=I_{N}.$ When $N$ is a power of two (i.e. $N=2^{m}$), we compute the transform of an $N\times C$ token matrix $X$ in $m=\log_{2}(N)$ iterative stages. We begin with ${\boldsymbol{W}}^{(0)}=X,$ and for each stage $s=1,\dots,m$, update
@@
    After $m$ stages, the result ${\boldsymbol{W}}^{(m)}$ equals $W\,X$. This structured, deterministic mixing redistributes each token’s information uniformly across all others while preserving every token’s $\ell_{2}$-norm.
  • Haar-Random Orthogonal Perturbation. At each denoiser block $k$ and diffusion time step $i$, we generate a dense orthogonal matrix ${\boldsymbol{Q}}_{k,i}\;\in\;\mathbb{R}^{N\times N}$ by first sampling ${\boldsymbol{A}}\sim\mathcal{N}(0,1)^{\,N\times N}$ and then computing its QR decomposition ${\boldsymbol{A}}={\boldsymbol{Q}}\,{\boldsymbol{R}}.$ We set ${\boldsymbol{Q}}_{k,i}={\boldsymbol{Q}}$. Since the entries of ${\boldsymbol{A}}$ are i.i.d. Gaussian, the orthogonal factor ${\boldsymbol{Q}}$ is distributed uniformly (with respect to the Haar measure) over the orthogonal group $\mathcal{O}(N)$, and by construction
    ${\boldsymbol{Q}}_{k,i}^{\!\top}{\boldsymbol{Q}}_{k,i}=I.$
    This yields an isotropic rotation in the $N$-dimensional token-index space, mixing all token positions globally without changing their $\ell_{2}$-norm.
Each method preserves the overall norm and energy of the embeddings but changes their local structure in uniquely effective ways. Empirically, these operations break up small-scale noise patterns that the denoiser might overfit, while still carrying global structure for high-quality sample generation. Among these methods, token shuffling typically provides the best overall performance: it is straightforward to implement, has minimal computational overhead, and consistently achieves significant improvements in both diversity and fidelity of generated samples.
Appendix B Societal impact

### v2

@@ Abstract
[…]
@@ 1 Introduction
Diffusion models ho2020denoising; song2020denoising; song2020score have emerged as the main methodology behind many successful generative models for images esser2024scaling; podell2023sdxl, videos blattmann2023stable; bar2024lumiere; yang2024cogvideox; yu2024efficient; tian2024videotetris, audio tian2025audiox; xue2024auffusion, and 3D objects cui2024neusdfusion; liu2024pi3d; yan2024dreamdissector. Despite their theoretical capacity to produce high-fidelity data, unguided diffusion models often suffer from poor sample quality, manifesting as visual artifacts, lack of semantic consistency, and insufficient sharp details (karras2024guiding). To mitigate these issues, classifier-free guidance (CFG) ho2022classifier has become the de facto approach to steer the generation process toward higher quality and more semantically aligned outputs. However, CFG is inherently limited to conditional generation and requires a specific training strategy that randomly replaces the input condition with a null condition.
@@
In this paper, we revisit existing attention-based guidance techniques and aim to bridge the gap between the effectiveness of CFG and that of training-free, condition-agnostic methods. We observe that CFG effectively recovers global structure and coarse details during the early denoising steps, whereas existing training-free methods tend to produce over-smoothed results at the same stage (see highlighted regions in Figure˜1). This is problematic because early denoising steps are critical for both image quality and prompt alignment, as they establish global structure, major shapes, and coarse semantics before the network begins refining fine details wang2023diffusion; dhariwal2021diffusion. If the model fails to capture correct semantics, object placement, or overall composition at this stage, it may never fully recover from that high-level mismatch in later refinements. This lack of sufficient early-step guidance likely explains why existing methods often yield only marginal improvements in prompt alignment and generation quality compared to CFG.
@@
Motivated by these insights, we introduce Token Perturbation Guidance (TPG), a novel method that directly perturbs intermediate token representations within the diffusion network, without requiring additional training or architectural changes. TPG employs token shuffling as the core operation to provide effective guidance signals. Specifically, token shuffling is (i) linear, (ii) preserves token norms, and (iii) disrupts local structure while maintaining global statistics. As shown in Figure˜1, TPG exhibits behavior similar to CFG and, compared to other training-free methods, more faithfully recovers both global structure and fine details at early denoising stages.
[…]
@@ 2 Related work
Score‑based diffusion models sohl2015deep; ho2020denoising; song2019generative; song2020score reverse a forward noising process by learning the score function, the gradient of the log data density, at multiple noise levels to progressively transform pure Gaussian noise into realistic samples (song2019generative). This principled estimator of the data distribution has outperformed previous generative modeling methods in both fidelity and mode coverage dhariwal2021diffusion. Since the score represents an explicit gradient field, the sampling trajectory can be guided by incorporating auxiliary gradients, leading to powerful guidance techniques such as classifier guidance (CG) dhariwal2021diffusion and classifier‑free guidance (CFG) ho2022classifier. These guidance methods significantly enhance image quality and prompt alignment, albeit at the cost of oversaturation (sadat2025eliminating) and reduced diversity (ho2022classifier; sadat2024cads).
@@ 3 Background
where $\nabla_{\mathbf{x}}\log p_{t}({{\mathbf{x}}})$ is the score function, representing the gradient of the log-density of the noisy data at time $t$. This gradient indicates the direction in which a sample should be updated to increase its likelihood w.r.t. the noisy data distribution. The unknown score function is typically approximated by a neural network $\mathbf{s}_{\theta}(\mathbf{x},t)$, trained using denoising score matching vincent2011connection. Given a noisy sample ${\mathbf{x}}_{t}$ at time $t$, the network predicts the score of the marginal distribution $p_{t}$. Training is performed by minimizing a weighted denoising score matching loss:
@@ 4 Token perturbation guidance
[…]
@@ 5 Comparing the behavior of TPG with other guidance methods
As shown in Figure˜2, TPG and CFG produce guidance vectors that are nearly orthogonal to the ground-truth noise throughout the trajectory, as indicated by cosine values close to zero. This aligns with observation shown in sadat2024eliminating, which demonstrates that the CFG update term can be decomposed into parallel and orthogonal components. Importantly, the parallel component primarily contributes to oversaturation. It is believed that orthogonality enables more effective steering of the predicted direction, and generally, the orthogonal component contributes to improved image quality. In contrast, PAG and SEG exhibit strong negative alignment during the middle steps, suggesting that they temporarily oppose the denoising direction. Figure˜2 (b) further shows that, compared to SEG and PAG, TPG more closely mirrors the behavior of CFG in terms of alignment between the predicted and ground-truth noise. Additionally, Figure˜2 (c) shows that TPG and CFG exhibit similar guidance magnitudes across all denoising steps, while SEG and PAG show substantially lower norms. This indicates that the update terms in CFG and TPG are more influential throughout the sampling process compared to that in PAG and SEG.
Figure˜3 complements our step-wise analysis by illustrating how each guidance method behaves in the frequency domain. TPG and CFG remain almost perfectly orthogonal to the ground-truth noise across all frequencies and time steps, except for a slight positive tilt in the lowest frequency bands. SEG, however, exhibits a clear negative stripe at medium frequencies during intermediate steps, confirming that its correction momentarily opposes the desired direction. The norm heatmaps further reveal that CFG and TPG inject a strong low-frequency signal in the early steps, while high-frequency modifications primarily occur in the later denoising steps. In contrast, SEG operates with less energy and displays a markedly different norm pattern when modifying high-frequency content, indicating a relatively weaker approach to detail refinement. This behavior aligns with Figure˜1, where the reduced energy across frequency bands leads to overly smooth generations in the initial steps.
@@ 6 Experiments
All experiments are conducted via official implementations and pre-trained checkpoints provided by publicly available repositories. We build upon the current open-source state-of-the-art Stable Diffusion XL (SDXL) podell2023sdxl as our primary baseline, and include the Stable Diffusion 2.1 (SD 2.1) CompVis2022StableDiffusionv12 to show the generality of our method and analysis. Moreover, TPG is easy to implement, and it can be added into existing diffusion models as a plug-and-play module with just a few lines of additional code. In all experiments, the compared methods are evaluated using their original configurations and default guidance scales. The guidance scale for TPG is fixed at 3.0.
@@ 6.1 Quantitative results
Table˜1 presents a quantitative comparison of TPG against vanilla SDXL, PAG, SEG, and CFG for both unconditional and conditional image generation tasks. In unconditional generation, TPG outperforms all compared methods by achieving notably lower FID and sFID scores (69.31 and 44.18, respectively), indicating better overall image quality and diversity. Additionally, TPG achieves the highest Inception Score (17.99) while maintaining a competitive Aesthetic Score, further reflecting the perceptual quality of TPG outputs. In conditional generation, CFG achieves the best results overall, with TPG closely following and consistently outperforming vanilla SDXL, PAG, and SEG in FID, sFID, and CLIP Score metrics.
Table˜2 further compares TPG against vanilla Stable Diffusion 2.1, PAG, and SEG in unconditional image generation. TPG consistently achieves the best performance, significantly outperforming the other methods with the lowest FID score (16.69) and highest Inception Score (36.28). It also maintains competitive results in terms of Aesthetic Score (5.97) and CLIP Score (29.30), highlighting TPG’s strength in producing high-quality and semantically aligned images.
@@ 6.2 Qualitative results
Figure˜4 highlights the differences in unconditional generations produced by various guidance methods. Vanilla SDXL, PAG, and SEG often generate abstract or repetitive textures lacking clear semantic structure, while TPG consistently produces well-structured and realistic scenes. Across various initial seeds, TPG is less prone to generating abstract patterns and more likely to form coherent spatial layouts with identifiable objects. Moreover, visual artifacts such as unnatural textures or distortions are more noticeable in outputs from baseline methods, whereas TPG reduces such artifacts, resulting in cleaner and more realistic generations.
We also show a qualitative comparison of conditional image generation across different guidance methods in Figure˜5. As can be seen, CFG and TPG consistently produce the highest-quality images, with sharp details and strong alignment to the text prompts. In contrast, other baselines such as SEG and PAG often generate outputs that deviate from the given condition, leading to less faithful semantic content. Furthermore, visual artifacts, such as distorted shapes or inconsistent textures, are more frequently observed in SEG and PAG outputs, while TPG exhibits improved robustness, generating cleaner and more semantically accurate images with fewer artifacts.
@@
Table 2: Quantitative comparison of TPG with vanilla Stable Diffusion 2.1 rombach2022high, PAG ahn2024self, and SEG hong2024smoothed for unconditional generation. TPG outperforms other baselines in all evaluated metrics.
@@
$\textrm{Inception Score}\uparrow$ 24.59 28.80 25.15 36.28
$\textrm{Aesthetic Score}\uparrow$ 5.07 5.93 5.83 5.97
$\textrm{Clip Score}\uparrow$ 27.74 29.03 28.53 29.30
@@
6.3 Ablating other norm-preserving perturbation methods
To assess the importance of shuffling during guidance, Table˜3 reports the results of various norm-preserving perturbation methods compared to the vanilla baseline. In addition to the shuffling strategy used in TPG, we evaluated Sign Flip, Hadamard, and Haar transforms, all described in Appendix˜A. While these three alternatives provide slight improvements in generation quality over the baseline, their gains are modest. In contrast, shuffling yields a substantially larger improvement in both FID and Inception Score, indicating significantly better image quality and diversity. Although all perturbations are orthogonal and preserve token norms, they influence the features differently. Shuffling randomly reorders tokens, disrupting local patterns while preserving recoverable global structure, which facilitates stronger guidance during inference. By comparison, Hadamard and Haar transformations mix all tokens together, potentially distorting useful information and weakening the guidance signal. Sign Flip merely alters the sign of each token, which may not offer a strong enough signal to steer samples effectively toward desirable regions of the data distribution.
@@ 7 Conclusion and discussion
Appendix
This appendix provides supplementary material to support the main paper.
@@ Appendix A Orthogonal token perturbation matrix designs
TPG directly applies structured perturbations to the intermediate token embeddings within the denoiser during inference. Specifically, consider the intermediate hidden‐state activations of the denoiser at a given layer, represented as a tensor ${\boldsymbol{H}}\in\mathbb{R}^{B\times N\times C}$, where $B$ denotes the batch size, $N$ the number of tokens, and $C$ the dimension of each token’s feature vector. At each denoiser layer $k$ and diffusion time step $i$, we apply the perturbation by multiplying the token-embedding with an orthogonal (or approximately orthogonal) matrix ${\boldsymbol{P}}_{k,i}\;\in\;\mathbb{R}^{N\times N}$ along the token dimension:
@@
  • Walsh–Hadamard Transform (WHT). The WHT uses a normalized Hadamard matrix ${\boldsymbol{W}}\in\mathbb{R}^{N\times N}$, whose entries are $\pm 1/\sqrt{N}$ and which satisfies ${\boldsymbol{W}}^{\top}{\boldsymbol{W}}=I_{N}.$ When $N$ is a power of two (i.e. $N=2^{m}$), we compute the transform of an $N\times C$ token matrix $X$ in $m=\log_{2}(N)$ iterative stages. We begin with ${\boldsymbol{W}}^{(0)}=X,$ and for each stage $s=1,\dots,m$, update
    $\begin{aligned} {\boldsymbol{W}}^{(s)}_{j,:}&={\boldsymbol{W}}^{(s-1)}_{j,:}+{\boldsymbol{W}}^{(s-1)}_{j+2^{s-1},:},\\ {\boldsymbol{W}}^{(s)}_{j+2^{s-1},:}&={\boldsymbol{W}}^{(s-1)}_{j,:}-{\boldsymbol{W}}^{(s-1)}_{j+2^{s-1},:},\end{aligned}\quad j=1,3,\dots,N-2^{s-1}+1.$
    After $m$ stages, the result ${\boldsymbol{W}}^{(m)}$ equals WXW\,X. This structured, deterministic mixing redistributes each token’s information uniformly across all others while preserving every token’s $\ell_{2}$-norm.
  • Haar‐Random Orthogonal Perturbation. At each denoiser block $k$ and diffusion time step $i$, we generate a dense orthogonal matrix ${\boldsymbol{Q}}_{k,i}\;\in\;\mathbb{R}^{N\times N}$ by first sampling ${\boldsymbol{A}}\sim\mathcal{N}(0,1)^{\,N\times N}$ and then computing its QR decomposition ${\boldsymbol{A}}={\boldsymbol{Q}}\,{\boldsymbol{R}}.$ We set ${\boldsymbol{Q}}_{k,i}={\boldsymbol{Q}}$. Since the entries of $\boldsymbol{A}$} are i.i.d. Gaussian, the orthogonal factor $\boldsymbol{Q}$} is distributed uniformly (with respect to the Haar measure) over the orthogonal group $\mathcal{O}(N)$, and by construction
    ${\boldsymbol{Q}}_{k,i}^{\!\top}{\boldsymbol{Q}}_{k,i}=I.$
    This yields an isotropic rotation in the $N$-dimensional token‐index space, mixing all token positions globally without changing their $\ell_{2}$-norm.
Each method preserves the overall norm and energy of the embeddings but changes their local structure in uniquely effective ways. Empirically, these operations break up small‐scale noise patterns that the denoiser might overfit, while still carrying global structure for high‐quality sample generation. Among these methods, token shuffling typically provides the best overall performance: it is straightforward to implement, has minimal computational overhead, and consistently achieves significant improvements in both diversity and fidelity of generated samples.
Table 4: Effect of guidance scale $\sigma$ on generative performance. Increasing $\sigma$ improves both FID and Inception Score up to $\sigma{=}3$, after which performance degrades, suggesting an optimal balance between diversity and fidelity at moderate noise levels.
@@
0 136.01 86.42 7.48 0.21 0.31
1 84.05 68.35 16.05 0.38 0.37
2 77.25 70.81 17.39 0.45 0.35
3 76.00 75.92 18.47 0.45 0.37
4 77.62 80.29 17.87 0.44 0.30
5 79.44 83.23 17.11 0.43 0.32
6 81.52 86.63 16.39 0.41 0.30
7 84.17 90.15 15.39 0.40 0.28
Table 5: Quantitative comparison of Vanilla SD3 esser2024scaling, TPG, and PAG ahn2024self for unconditional generation, evaluated over 5k generated samples. TPG achieves the lowest FID and highest Inception Score, showing improved fidelity and diversity over other baselines.
[…]
Vanilla SD3 esser2024scaling 113.86 91.09 11.06 0.26 0.28
PAG ahn2024self 138.08 216.65 9.13 0.25 0.15
TPG (Ours) 83.01 71.59 13.34 0.46 0.42
Table 6: Quantitative comparison of Vanilla SD3 esser2024scaling, CFG ho2022classifier, TPG, and PAG ahn2024self for conditional generation, evaluated over 5k generated samples. CFG ho2022classifier achieves the highest precision, while TPG attains a balanced trade-off between quality and diversity.
[…]
Vanilla SD3 esser2024scaling 33.81 63.17 29.62 0.45 0.45
CFG ho2022classifier 21.22 58.14 42.10 0.70 0.42
PAG ahn2024self 41.32 117.03 25.65 0.42 0.30
TPG (Ours) 21.37 57.01 34.49 0.56 0.49
Appendix B More ablation
Effect of guidance scale on TPG
We investigate the impact of the guidance scale $\sigma$ on the quality and diversity of generated samples. As shown in Table˜4, increasing $\sigma$ from 0 to 3 consistently improves both FID and Inception Score, indicating better fidelity and diversity. The performance peaks at $\sigma{=}3$ with the lowest FID (76.00) and the highest Inception Score (18.47), suggesting an optimal trade-off between structure preservation and generative diversity. Beyond this point, larger guidance scales ($\sigma>3$) lead to gradual degradation across all metrics, implying that excessive stochasticity introduces artifacts and reduces consistency in generation.
Compatibility with ViT-based models
Unlike U-Net architectures that feature explicit multi-resolution hierarchies, DiT (Transformer-based diffusion models) lacks a coarse-to-fine synthesis structure. Instead, it distributes functionality more uniformly across layers, making it less clear which layers capture high-level semantic features avrahami2025stable. This structural difference can complicate the use of traditional image editing and perturbation methods. Furthermore, residual connections in U-Net architectures help the model recover degraded latent tokens that have been intentionally perturbed. In contrast, DiT architectures consist of consecutive Transformer layers without residual connections between them. This design causes degradation to accumulate across layers, leading to significant signal loss or distortion. As a result, the guidance signal becomes less effective, often pushing the generated images out of distribution. Consequently, conventional perturbation-based methods are not directly applicable to DiT architectures, as demonstrated in Tables˜5 and 6, where PAG ahn2024self performs even worse than the vanilla SD3 model esser2024scaling.
However, this limitation does not apply to our proposed TPG method. By leveraging the linear property of the shuffling operation, we can revert shuffled tokens to their original positions, effectively restoring the degraded latent tokens and enabling more meaningful guidance signals. Specifically, for DiT-based architectures such as SD3 esser2024scaling, we shuffle only a small fraction of input tokens before each Transformer block and immediately unshuffle them afterward, returning the permuted tokens to their original locations. Through this unshuffling process, TPG enables perturbation-based guidance to performe effectively within DiT architectures.
To evaluate the generalization ability of TPG beyond the U-Net architecture, we compare it with existing perturbation-based guidance methods, including PAG ahn2024self, using the Stable Diffusion 3 model. The results for both unconditional and conditional generation are shown in Tables˜5 and 6. Unlike prior perturbation-based approaches that were specifically designed for U-Net and fail to transfer effectively, TPG demonstrates strong cross-architecture compatibility. In the unconditional setting, TPG achieves the best overall performance, significantly reducing FID and improving Inception Score compared to both Vanilla SD3 esser2024scaling and PAG ahn2024self. In the conditional setting, TPG performs competitively with CFG ho2022classifier, achieving comparable FID and recall while maintaining high visual fidelity. These results confirm that TPG generalizes effectively across different architectures and retains its effectiveness without modification.
Ablation on layer selection
Following prior works such as SEG hong2024smoothed and PAG ahn2024self for U-Net-based models, we examined perturbing the downsampling, mid, and upsampling layers separately. Table˜7 shows that TPG is most effective when the perturbation is applied only to the downsampling layers (i.e., the encoder part of the U-Net). We also experimented with combinations of down, mid, and up layers, but these did not lead to improvements and, in some cases, resulted in degraded performance.
Table 7: Quantitative comparison of shuffling applied to down, mid, and up layers for unconditional generation using SDXL, evaluated over 5k generated samples.
[…]
Down layers 76.00 75.92 18.47 0.45 0.37
Mid layers 100.80 142.40 12.27 0.29 0.32
Up layers 102.28 154.87 12.81 0.30 0.23
Exploring non-norm preserving perturbations
Table 8: Quantitative comparison of Token Shuffling (ours), Token Blurring, and Vanilla for unconditional generation using SDXL, evaluated over 5k generated samples. Token Shuffling substantially improves all quality metrics, demonstrating its effectiveness in enhancing generative fidelity and diversity.
[…]
Vanilla 136.01 86.42 7.48 0.21 0.31
Token Blurring 157.67 184.40 6.70 0.18 0.23
Token Shuffling (ours) 76.00 75.92 18.47 0.45 0.37
We further investigate the effect of non–norm-preserving perturbations by introducing Token Blurring, which applies a Gaussian blur kernel to the input tokens. As shown in Table˜8, Token Blurring leads to a clear degradation across all evaluation metrics, with higher FID and lower Inception Score compared to our Token Shuffling strategy. In contrast, Token Shuffling maintains feature magnitude consistency while introducing structured diversity, resulting in improved fidelity and perceptual quality. These results indicate that preserving the token norm during perturbation is crucial for stable and semantically coherent generation.
Appendix C Limitations and future work
Despite these strengths, as with CFG ho2022classifier itself, TPG still requires two forward passes through the diffusion network, leading to increased sampling time compared to the unguided case. Additionally, although TPG significantly improves quality in most situations, the guidance term may remain limited in extreme out-of-distribution scenarios that are not captured by the learned distribution of the base model. We consider addressing these limitations an interesting direction for future research.
Appendix D Societal impact